OpenStudy (anonymous):

I need to write the slope-intercept equation for a line that passes through the points (-3,0) and (3,1). y=mx+b, so using y1-y2/x1-x2 I get a slope of 1/6, changing the equation into y=1/6x+b How do I find b?

OpenStudy (amistre64):

or to relate it to finding slope ... we can progress thru these steps \[\frac{y-Py}{x -Px} = m\] \[y-Py = m(x -Px)\] \[y-Py = mx -mPx\] \[y = mx -mPx+Py\] \[y = mx +[-mPx+Py]\] \[y = mx +b\]

OpenStudy (anonymous):

Can Py be any of the y variables in the two given points? Likewise for Px?

OpenStudy (amistre64):

in a sense, yes; but make sure your Py and Px are from the same given point :)

OpenStudy (amistre64):

once you know: y=mx+b, you can actually find "b" by using either given point as well. Py = m(Px) + b Py - mPx = b; which amounts to what i wrote above i believe